OpenStudy (anonymous):

Eric knew his trip to Lawrenceburg was 548 miles. If he drove for 8.9 hours, what mph did he average on his trip? You may use the formula d = rt. Round your answer to the nearest whole number. A. 60 mph B. 61 mph C. 62 mph Eric knew his trip to Lawrenceburg was 548 miles. If he drove for 8.9 hours, what mph did he average on his trip? You may use the formula d = rt. Round your answer to the nearest whole number. A. 60 mph B. 61 mph C. 62 mph @Mathematics

OpenStudy (turingtest):

d=vt v=d/t=548/8.9=61.573... which rounds to 62mph